# RealTraj: Towards Real-World Pedestrian Trajectory Forecasting

### v2

@@ Abstract
This paper jointly addresses three key limitations in conventional pedestrian trajectory forecasting: pedestrian perception errors, real-world data collection costs, and person ID annotation costs. We propose a novel framework, RealTraj, that enhances the real-world applicability of trajectory forecasting. Our approach includes two training phases—self-supervised pretraining on synthetic data and weakly-supervised fine-tuning with limited real-world data—to minimize data collection efforts. To improve robustness to real-world errors, we focus on both model design and training objectives. Specifically, we present Det2TrajFormer, a trajectory forecasting model that remains invariant in tracking noise by using past detections as inputs. Additionally, we pretrain the model using multiple pretext tasks, which enhance robustness and improve forecasting performance based solely on detection data. Unlike previous trajectory forecasting methods, our approach fine-tunes the model using only ground-truth detections, significantly reducing the need for costly person ID annotations. In the experiments, we comprehensively verify the effectiveness of the proposed method against the limitations, and the method outperforms state-of-the-art trajectory forecasting methods on multiple datasets. The code will be released at project page.
@@ 1 Introduction
The pedestrian trajectory forecasting task aims to predict the future positions of pedestrians based on their past observed states, a critical capability for understanding motion and behavioral patterns [49]. This task plays a crucial role in real-world applications such as social robotics navigation [28, 17], autonomous driving [10], and surveillance systems [14]. Over the years, deep learning models have dominated this domain, with the RNN-based [1, 24, 32], the graph-based [2, 52, 3, 65], and the Transformer-based approaches [22, 21, 53, 50] setting new performance benchmarks on large-scale datasets [76, 45, 33, 48, 4, 71, 40]. However, most existing approaches are evaluated under fixed, well-known benchmarks with ideal experimental conditions, such as perfectly annotated trajectories on the pre-recorded videos. This limits their applicability to real-world scenarios, necessitating further performance enhancements for practical deployment. In this paper, we address three key limitations present in existing approaches, as described below.
@@
Real-World Data Collection Costs. Conventional trajectory forecasting approaches are typically trained on large-scale datasets and evaluated on domains with data similar to the training conditions (e.g., consistent camera angles and sensor setups). This often overlooks the substantial costs associated with real-world data collection, which often requires significant manual effort. Thus, it is crucial to develop prediction models that can be trained effectively while minimizing real-world data collection efforts.
@@
Person ID Annotation Costs. In addition to the challenge of large-scale real-world data collection, most existing trajectory forecasting models rely on fully supervised training methods. This approach necessitates ground-truth data for future trajectories, including precise pedestrian positions and consistent person identifiers across frames, to minimize the loss between predicted and actual trajectories.
@@
In this paper, we propose a novel trajectory forecasting framework, RealTraj, designed to enhance the real-world applicability of trajectory forecasting by addressing three key limitations. Our approach consists of two training phases—self-supervised pretraining and weakly-supervised fine-tuning—followed by an inference phase. First, we leverage large-scale synthetic trajectory data generated from a simulator [58] to effectively learn trajectory patterns using only synthetic data, reducing the need for extensive real-world data collection (addressing the second limitation). However, relying solely on clean synthetic and real-world data can lead to reduced robustness against real-world errors. To mitigate this, we enhance both the model design and training objectives: we introduce Det2TrajFormer, a trajectory forecasting model designed to be invariant to tracking noise by utilizing past detections as inputs. Additionally, we pretrain the model with multiple self-supervised tasks to improve robustness against perception errors (addressing the first limitation) and enhance forecasting performance solely from detections. Unlike previous trajectory forecasting approaches, we propose training the model exclusively on ground-truth detections during the fine-tuning phase, thereby reducing person ID annotation costs (addressing the third limitation). We further introduce an acceleration regularization term to discourage abrupt changes in acceleration, resulting in smoother and more realistic trajectory forecastings. During inference, Det2TrajFormer takes detection results as input to forecast future trajectories.
@@
In summary, our contributions are threefold: (1) We introduce a novel trajectory forecasting framework with self-supervised pretraining on large-scale synthetic data and weakly-supervised fine-tuning on limited real-world data, enhancing real-world applicability by addressing pedestrian perception errors, data collection costs, and person ID annotation costs. (2) We propose three self-supervised tasks to improve robustness against pedestrian perception errors and enhance trajectory forecasting accuracy using only detections. (3) We reduce person ID annotation costs by training the model solely on ground-truth detections via proposed weakly-supervised loss during the fine-tuning phase.
@@ 2.1 Robust Trajectory Forecasting
Pedestrian trajectory forecasting models aim to predict future positions based on observed trajectories, with deep learning methods demonstrating strong performance due to their representational capabilities [1, 24, 41, 62, 31, 37, 22, 51, 26, 36, 52, 23, 64, 38, 50, 65, 21]. Various architectural designs, such as the RNN-based [1, 24, 32], the graph-based [2, 52, 3, 65], and more recent Transformer-based approaches [22, 21, 53, 50], have been explored to improve forecasting accuracy.
Despite the significant advancements, most trajectory forecasting models operate under idealized conditions, relying on ground-truth past trajectories for training and evaluation while ignoring the impact of imperfect inputs. Recent studies have increasingly recognized the challenges posed by perception module errors [61, 18, 68, 74, 72, 54, 60, 69, 66, 9, 16] and adversarial attacks [73, 5, 75, 29]. This work focuses on enhancing robustness against perception module errors, which are unavoidable in real-world deployments and can significantly degrade prediction accuracy [70, 61, 60, 69, 19]. While some approaches address incomplete observations due to occlusions or limited viewpoints (i.e., detection errors) [18, 68, 72, 54, 66, 9], others tackle tracking errors [70, 61, 60, 74, 16]. However, these methods often utilize architectures and objectives tailored to specific error types, limiting their applicability in diverse real-world scenarios. In this work, we aim to develop a robust model capable of handling both detection and tracking errors.
@@ 2.2 Adapting Trajectory Forecasting
Conventional trajectory forecasting models often depend heavily on specific training data domains, overlooking the substantial costs and manual effort required for real-world data collection. To address this challenge, recent research has focused on lightweight methods for adapting pretrained predictors to new data. Some approaches target cross-domain transfer [67, 27, 35], while others emphasize online adaptation [44, 34, 8], continual learning [39, 64, 55], and prompting-based strategies [57]. These methods have demonstrated notable improvements in forecasting performance. However, they often assume access to ground-truth trajectories, which require costly person ID annotations. To reduce this burden, we propose fine-tuning models using only ground-truth detections, thereby eliminating the extensive need for person ID annotations.
@@ 3.1 Problem Formulation
We aim to forecast the future trajectory of a target pedestrian in the scene based on detections from observed frames, including the non-targeted pedestrians in the environment, from observed frames. Unlike prior work, our approach skips the tracking step and instead uses detections directly, thereby avoiding the propagation of tracking errors into the forecasting task. Formally, let $Y=(y_{1},y_{2},\dots,y_{T_{pred}})\in\mathbb{R}^{T_{pred}\times 2}$ denote the future trajectory of the target pedestrian over $T_{pred}$ time steps where each position is represented as a tuple $y_{t}=(u_{t},v_{t})\in\mathbb{R}^{2}$, and let $X=(X_{1},X_{2},\dots,X_{T_{obs}})\in\mathbb{R}^{KT_{obs}\times 2}$ represent the set of past detections of $K$ pedestrians over $T_{obs}$ time steps. The detections at time $t$ are defined as $X_{t}=(x_{t}^{1},x_{t}^{2},\dots,x_{t}^{K})\in\mathbb{R}^{K\times 2}$, where each detection is the position of pedestrian $x_{t}^{k}=(u_{t}^{k},v_{t}^{k})\in\mathbb{R}^{2}$ at timestep $t$ [^1] Note that, unlike the conventional trajectory forecasting approaches, our method takes a set of detections as input at each frame. Therefore, it is not guaranteed that $x_{t}^{1}$ and $x_{t+1}^{1}$ correspond to detections of the same pedestrian.. The target pedestrian is one of the $K$ pedestrians in the last observed frame ($t=T_{obs}$). Our goal is to learn a function $\mathcal{F}(X)$ that maps the input detections $X$ to a predicted future trajectory of a target pedestrian $Y$, such that $Y=\mathcal{F}(X)$. In this work, we introduce a novel weakly-supervised trajectory forecasting setting. Large-scale synthetic trajectory data are available during pretraining, and real-world limited detection data are available during fine-tuning.
@@ 3.2 Overview
Our framework consists of two training phases—self-supervised pretraining and weakly-supervised fine-tuning—along with an inference phase, as illustrated in Fig. 2. We leverage large-scale synthetic trajectory data from a trajectory simulator [58] to learn effective trajectory patterns, reducing the need for extensive real-world data collection. To ensure real-world robustness, we approach from both the training scheme and the proposed trajectory forecasting model $\mathcal{F}$, Det2TrajFormer. During the pretraining phase, we leverage multiple pretext tasks to improve robustness and forecasting accuracy using large-scale synthetic trajectory data. In the fine-tuning phase, training is conducted solely on ground-truth detections, eliminating the need for costly person ID annotations. During inference, Det2TrajFormer predicts future trajectories based on detection inputs.
@@
$\begin{split}[\hat{H},\hat{Q}]&=\mathcal{E}(\mathcal{G}(X),Q),\\ \hat{Y}&=\mathcal{P}_{f}(\hat{Q}).\end{split}$ (1)
@@ 3.3 Self-Supervised Pretraining Phase
Using synthetically generated trajectories, we train the model on the primary forecasting task while incorporating pretext tasks to enhance performance and improve robustness against detection errors. From the synthetically generated trajectories, we generate a set of past detections $X$ by removing the identity information. Additionally, we simulate the miss-detection and localization errors in the inputs by randomly masking and adding Gaussian noises to $X$, resulting in the corrupted past detections, $X^{C}$. We propose three pretext tasks that aim to reconstruct the removed person ID information from $X$ (person ID reconstruction), reconstruct the masked detections from $X^{C}$ (unmasking), and denoise the $X^{C}$ (denoising). During the pretraining, we additionally employ three heads $\mathcal{P}_{R}\in\mathbb{R}^{d\rightarrow d_{I}}$, $\mathcal{P}_{M}\in\mathbb{R}^{d\rightarrow 2}$, and $\mathcal{P}_{D}\in\mathbb{R}^{d\rightarrow 2}$, where $d$ and $d_{I}$ denote the feature dimension and dimension of person ID embedding.
@@
Unmasking Task. The goal of the unmasking pretext task is to encourage the encoder to learn robust features against miss-detections. Specifically, we input $\hat{H}$ obtained from corrupted detection $X^{C}$ by Eq. 1 into the unmasking head $\mathcal{P}_{M}$, as follows:
$\hat{X}_{unmask}=\mathcal{P}_{M}(\hat{H}).$ (2)
Denoising Task. The goal of the denoising pretext task is to encourage the encoder to learn robust features that mitigate localization errors. Similarly, the denoising is formulated as follows:
$\hat{X}_{denoise}=\mathcal{P}_{D}(\hat{H}).$ (3)
Person Identity Reconstruction Task.
The person identity reconstruction pretext task aims to enhance the model’s ability to effectively associate pedestrian instances across observation frames. To achieve this, the network is trained to reconstruct pedestrian identities from detections $X$ by predicting the person identity embeddings. The person identity reconstruction task is formulated as follows:
$\hat{I}=\mathcal{P}_{R}(\hat{H}),$ (4)
@@
Loss.
The model is trained to jointly minimize the losses for the four tasks mentioned above:
@@
where $\mathcal{L}{{}_{F}}$, $\mathcal{L}_{M}$, $\mathcal{L}_{D}$, and $\mathcal{L}_{I}$ represent the trajectory forecasting, unmasking, denoising, and person identity reconstruction losses, respectively. We use the mean square error (MSE) as the loss function between the targets and predictions for each task. The terms $\alpha$, $\beta$, and $\gamma$ are weight hyperparameters.
@@ 3.4 Weakly-Supervised Fine-tuning Phase
We fine-tune the model in a weakly-supervised manner by using future ground-truth detections $X_{T_{obs}+1:T_{obs}+T_{pred}}$ instead of trajectories $Y$, thereby reducing ID annotation costs in real-world data. To achieve this, we propose to employ two losses. The first loss is calculated between the predicted future position $\hat{y_{t}}$ and the closest ground-truth detection $d_{t}^{c}$ at each future timestep, summing over the $T_{pred}$ timesteps as follows:
$d_{t}^{c}=\underset{k\in K}{\operatorname{argmin}}\|\hat{y}_{t}-d_{t}^{k}\|_{2},$ (6)
@@
$\mathcal{L}_{finetune}=\mathcal{L}_{W}+\lambda\mathcal{L}_{Reg},$ (9)
@@
Refer to caption Refer to caption
@@
Table 1: Comparisons with the current state-of-the-art method on JRDB, JTA, ETH-UCY, and SDD under few-shot evaluation. We summarize the ADE metric with different percentages of labeled data. The unit for ADE is meters for JRDB, JTA, and ETH-UCY, and pixels for SDD. Note that the upper bounds for $100\%$ supervised-RealTraj are $0.35$, $0.92$, $0.51$, and $15.99$ for JRDB, JTA, ETH-UCY, and SDD, respectively. The best results are bolded, and the second-best results are underlined.
(a) JRDB
Setting Method $0.1\%$ $0.5\%$ $1\%$ $2\%$ $5\%$
Supervised EqMotion [65]${}_{\text{CVPR'}23}$ 0.45 0.48 0.43 0.60 0.43
Social-Trans [50]${}_{\text{ICLR'}24}$ 0.68 0.48 0.47 0.43 0.43
RealTraj (Ours) 0.43 0.41 0.40 0.38 0.38
Weakly
Supervised
RealTraj (Ours) 0.43 0.41 0.40 0.39 0.38
(b) JTA
Setting Method $0.1\%$ $0.5\%$ $1\%$ $2\%$ $5\%$
Supervised EqMotion [65]${}_{\text{CVPR'}23}$ 1.76 1.68 1.55 1.55 1.55
Social-Trans [50]${}_{\text{ICLR'}24}$ 4.80 3.33 2.91 2.78 2.78
RealTraj (Ours) 1.38 1.25 1.22 1.20 1.07
Weakly
Supervised
RealTraj (Ours) 2.13 1.98 1.68 1.46 1.23
(c) ETH-UCY
Setting Method $0.1\%$ $0.5\%$ $1\%$ $2\%$ $5\%$
Supervised EqMotion [65]${}_{\text{CVPR'}23}$ 0.70 0.56 0.53 0.53 0.53
Social-Trans [50]${}_{\text{ICLR'}24}$ 1.05 0.62 0.55 0.52 0.53
RealTraj (Ours) 0.56 0.55 0.53 0.52 0.52
Weakly
Supervised
RealTraj (Ours) 0.56 0.56 0.56 0.53 0.53
(d) SDD
Setting Method $0.1\%$ $0.5\%$ $1\%$ $2\%$ $5\%$
Supervised EqMotion [65]${}_{\text{CVPR'}23}$ 24.67 19.95 18.53 19.03 17.61
Social-Trans [50]${}_{\text{ICLR'}24}$ 45.02 32.67 23.68 21.80 19.03
RealTraj (Ours) 22.76 20.16 19.22 18.49 17.03
Weakly
Supervised
RealTraj (Ours) 23.49 20.33 19.74 18.99 17.89

### v3

@@ Abstract
This paper jointly addresses three key limitations in conventional pedestrian trajectory forecasting: pedestrian perception errors, real-world data collection costs, and person ID annotation costs. We propose a novel framework, RealTraj, that enhances the real-world applicability of trajectory forecasting. Our approach includes two training phases—self-supervised pretraining on synthetic data and weakly-supervised fine-tuning with limited real-world data—to minimize data collection efforts. To improve robustness to real-world errors, we focus on both model design and training objectives. Specifically, we present Det2TrajFormer, a trajectory forecasting model that remains invariant to tracking noise by using past detections as inputs. Additionally, we pretrain the model using multiple pretext tasks, which enhance robustness and improve forecasting performance based solely on detection data. Unlike previous trajectory forecasting methods, our approach fine-tunes the model using only ground-truth detections, reducing the need for costly person ID annotations. In the experiments, we comprehensively verify the effectiveness of the proposed method against the limitations, and the method outperforms state-of-the-art trajectory forecasting methods on multiple datasets. The code will be released at project page.
@@ 1 Introduction
The pedestrian trajectory forecasting task aims to predict the future positions of pedestrians based on their past observed states, a critical capability for understanding motion and behavioral patterns [60]. This task plays a crucial role in real-world applications such as social robotics navigation [34, 22], autonomous driving [14], and surveillance systems [19]. Over the years, deep learning models have dominated this domain with the various model architectures, such as RNNs [1, 29, 39, 5], graphs [2, 63, 3, 79], Transformers [27, 26, 64, 61] and diffusions [48, 28, 6, 30], and these models are evaluated on large-scale benchmarks [90, 56, 40, 59, 7, 85, 50]. However, existing approaches are evaluated under fixed, well-known benchmarks with ideal experimental conditions, such as completely annotated trajectories with the pre-recorded videos. This limits their applicability to real-world scenarios, necessitating further performance enhancements for practical deployment. In this paper, we address three key limitations present in existing approaches, as described below (Fig. 1).
@@
Real-World Data Collection Costs. Conventional trajectory forecasting approaches are typically trained on large-scale datasets and evaluated on data domains similar to the training (e.g., consistent camera angles and sensor setups). However, these methods often overlook the substantial costs associated with real-world data collection, including the acquisition of raw sensor data such as point clouds or videos, which requires significant manual effort. Thus, it is crucial to develop prediction models that can be trained effectively while minimizing real-world data collection efforts.
@@
Person ID Annotation Costs. In addition to the challenge of large-scale real-world data collection, most existing trajectory forecasting models rely on fully supervised training. This approach necessitates ground-truth future trajectories, including precise pedestrian positions and consistent person identities across frames.
@@
In this paper, we propose a novel trajectory forecasting framework, RealTraj, designed to enhance the real-world applicability of trajectory forecasting by jointly addressing aforementioned three limitations. Our approach consists of two training phases—self-supervised pretraining and weakly-supervised fine-tuning—followed by an inference phase. First, we leverage large-scale synthetic trajectory data generated from a simulator [70] to effectively learn trajectory patterns using only synthetic data, reducing the need for extensive real-world data collection (addressing the second limitation). However, relying solely on clean synthetic data weakens robustness against real-world errors.
To mitigate this, we enhance both the model design and training objectives: we introduce Det2TrajFormer, a trajectory forecasting model designed to be invariant to tracking noise by only inputs past detections (without any person indices across frames). Additionally, inspired by the recent progress of self-supervised learning mechanism [31, 74, 78], we pretrain the model with multiple self-supervised pretext tasks, such as unmasking, denoising, and person identity reconstruction, to improve robustness against perception errors (addressing the first limitation) while enhancing forecasting performance solely from detections.
Finally, unlike previous trajectory forecasting approaches, we propose training the model exclusively on ground-truth detections during the fine-tuning phase, thereby reducing person ID annotation costs (addressing the third limitation) on the real data. We further introduce an acceleration regularization term to discourage abrupt changes in acceleration, resulting in smoother and more realistic trajectory forecastings. During inference, Det2TrajFormer takes detection results as input to forecast future trajectories.
@@
In summary, our contributions are three-fold: (1) We introduce a novel trajectory forecasting framework, RealTraj, with self-supervised pretraining on large-scale synthetic data and weakly-supervised fine-tuning on limited real-world data, enhancing real-world applicability by jointly addressing pedestrian perception errors, data collection costs, and person ID annotation costs within a unified and single framework. (2) We propose three self-supervised pretext tasks to improve robustness against pedestrian perception errors and enhance trajectory forecasting accuracy using only detections. (3) We reduce person ID annotation costs by training the model solely on ground-truth detections via proposed weakly-supervised loss.
@@ 2.1 Robust Trajectory Forecasting
Pedestrian trajectory forecasting models aim to predict future positions based on observed trajectories. Deep learning methods demonstrating strong performance due to their representational capabilities [1, 29, 51, 75, 38, 47, 27, 62, 32, 46, 63, 28, 77, 48, 61, 79, 26, 30, 15].
Despite the significant advancements, most trajectory forecasting models operate under idealized conditions, relying on ground-truth past trajectories for training and evaluation while ignoring the impact of imperfect inputs. Recent studies have increasingly recognized the challenges posed by perception module errors [73, 23, 82, 88, 86, 65, 72, 83, 80, 13, 21, 67, 42, 43] and adversarial attacks [87, 8, 89, 35]. This work focuses on enhancing robustness against perception module errors, which are unavoidable in real-world deployments and can significantly degrade prediction accuracy [84, 73, 72, 83, 24]. Some approaches address incomplete observations, encompassing missing data [23, 82, 65, 13], momentary visibility [67, 42, 52], variations in observation length [80, 43], and out-of-sight instances [86] caused by occlusions or limited viewpoints (i.e., detection errors). Others focus on mitigating tracking errors [84, 73, 72, 88, 21]. Unlike the previous approaches, which independently tailored the approach for the specific error types, we aim to improve the robustness jointly against both detection (miss and wrong detection) and tracking errors within a single framework.
@@ 2.2 Adapting Trajectory Forecasting
Conventional trajectory forecasting models often heavily depend on specific training data domains, overlooking the substantial costs and manual effort required for real-world data collection, such as collecting the videos to obtain the trajectories. To address this challenge, recent research has focused on lightweight methods for adapting pretrained prediction models to newly captured data. Some approaches target cross-domain transfer [81, 33, 44], while others emphasize online adaptation [55, 41, 12], continual learning [49, 77, 66], and prompting-based strategies [69]. These methods have demonstrated notable improvements in forecasting performance. However, they often assume access to ground-truth trajectories, which require costly person ID annotations. To reduce this burden, we propose to fine-tune the model using only ground-truth detections, thereby eliminating the extensive need for person ID annotations.
@@ 3.1 Problem Formulation
We aim to forecast the future trajectory of a target pedestrian in the scene based on detections from observed frames, including the non-targeted pedestrians in the environment. Unlike prior work, our approach skips the tracking step and instead uses detections directly, thereby avoiding the propagation of tracking errors into the forecasting task. Formally, let $Y=(y_{1},y_{2},\dots,y_{T_{pred}})\in\mathbb{R}^{T_{pred}\times 2}$ denote the future trajectory of the target pedestrian over $T_{pred}$ time steps where each position is represented as a tuple $y_{t}=(u_{t},v_{t})\in\mathbb{R}^{2}$, and let $X=(X_{1},X_{2},\dots,X_{T_{obs}})\in\mathbb{R}^{KT_{obs}\times 2}$ represent the set of past detections of $K$ pedestrians over $T_{obs}$ time steps. The detections at time $t$ are defined as $X_{t}=(x_{t}^{1},x_{t}^{2},\dots,x_{t}^{K})\in\mathbb{R}^{K\times 2}$, where each detection is the position of pedestrian $x_{t}^{k}=(u_{t}^{k},v_{t}^{k})\in\mathbb{R}^{2}$ at timestep $t$ [^1] Unlike the conventional trajectory forecasting approaches, our method takes a set of detections as input at each frame. Therefore, it is not guaranteed that $x_{t}^{1}$ and $x_{t+1}^{1}$ correspond to detections of the same pedestrian.. The target pedestrian is one of the $K$ pedestrians in the last observed frame ($t=T_{obs}$). The model can be applied $K$ times to predict $K$ pedestrians’ trajectories. The target pedestrian is specified by translating the detections to make one of the detections in the last frame to the origin. The goal is to learn a function $\mathcal{F}(X)$ that maps the input detections $X$ to a predicted future trajectory $Y$ of a target pedestrian, such that $Y=\mathcal{F}(X)$.
@@ 3.2 Overview
Our framework consists of two training phases—self-supervised pretraining and weakly-supervised fine-tuning—along with an inference phase, as illustrated in Fig. 2. We leverage large-scale synthetic trajectory data from a simulator [70] to effectively learn patterns while reducing the need for extensive real-world data collection. To ensure real-world robustness, we approach from both the training scheme and the trajectory forecasting model $\mathcal{F}$, Det2TrajFormer. During the pretraining, we leverage multiple pretext tasks to improve robustness and forecasting accuracy using large-scale synthetic trajectory data. During the fine-tuning, the model is fine-tuned solely on real-world ground-truth detections, eliminating the need for costly person ID annotations. During inference, Det2TrajFormer predicts future trajectories from observed detections.
@@
$\begin{split}[\hat{H},\hat{Q}]&=\mathcal{E}([\mathcal{G}(X),Q]),\\ \hat{Y}&=\mathcal{P}_{f}(\hat{Q}),\end{split}$ (1)
where []delimited-[][$\cdot$][ ⋅ ] denotes the concatenation operation along the token axis. Following recent works [25, 11], our approach can incorporate multiple forecasting heads ($\mathcal{P}_{f}^{n}$) to generate $N$ possible future predictions: $\hat{Y}^{n}=(\hat{y}_{1}^{n},\hat{y}_{2}^{n},\dots,\hat{y}_{T_{pred}}^{n})\in\mathbb{R}^{T_{pred}\times 2},\quad n=1,2,\dots,N$. For simplicity, we describe our method using a single future prediction setting.
@@ 3.3 Self-Supervised Pretraining Phase
Using synthetically generated trajectories, we train the model on the primary forecasting task while incorporating pretext tasks to enhance performance and improve robustness against detection errors. From the synthetically generated trajectories, we generate a set of past detections $X$ by removing the identity information. Additionally, we simulate the miss-detection and localization errors to $X$ by randomly masking and adding Gaussian noises, resulting in the corrupted past detections, $X^{C}$. We propose three pretext tasks that aim to reconstruct the masked detections from $X^{C}$ (unmasking), denoise the $X^{C}$ (denoising), and reconstruct the removed person ID information from $X$ (person ID reconstruction). Only during the pretraining, we additionally employ three heads $\mathcal{P}_{M}\in\mathbb{R}^{d\rightarrow 2}$, $\mathcal{P}_{D}\in\mathbb{R}^{d\rightarrow 2}$, and $\mathcal{P}_{R}\in\mathbb{R}^{d\rightarrow d_{I}}$, where $d_{I}$ denotes the dimension of person ID embedding.
@@
Unmasking Task. The goal of the unmasking pretext task is to encourage the encoder $\mathcal{E}$ to learn robust features against miss-detections. Specifically, we input $\hat{H}$ obtained from corrupted detection $X^{C}$ by Eq. 1 into the unmasking head $\mathcal{P}_{M}$, as follows: $\hat{X}_{unmask}=\mathcal{P}_{M}(\hat{H}).$
Denoising Task. The goal of the denoising pretext task is to encourage the encoder to learn robust features that mitigate localization errors. Similarly, the denoising is formulated as follows: $\hat{X}_{denoise}=\mathcal{P}_{D}(\hat{H}).$
Person Identity Reconstruction Task. The person identity reconstruction pretext task aims to enhance the model’s ability to effectively associate pedestrian instances across observation frames. To achieve this, the model is trained to reconstruct pedestrian identities from detections $X$ by predicting the person identity embeddings. The person identity reconstruction task is formulated as follows: $\hat{I}=\mathcal{P}_{R}(\hat{H}),$
@@
Loss. The model is trained to minimize the losses for the four tasks mentioned above:
@@
where $\mathcal{L}{{}_{F}}$, $\mathcal{L}_{M}$, $\mathcal{L}_{D}$, and $\mathcal{L}_{I}$ represent the trajectory forecasting [^2] For the multi-future prediction, we employ winner-take-all strategy which only optimizes the best prediction with minimal average prediction error to the ground truth, following  [11]., unmasking, denoising, and person identity reconstruction losses, respectively. We use the mean square error (MSE) loss between the targets and predictions for each task. The terms $\alpha$, $\beta$, and $\gamma$ are weight hyperparameters.
@@ 3.4 Weakly-Supervised Fine-tuning Phase
We fine-tune the model in a weakly-supervised manner by using future ground-truth detections $X_{T_{obs}+1:T_{obs}+T_{pred}}$, thereby reducing ID annotation costs in real-world data. To achieve this, we propose to employ two losses. The first loss is calculated between the predicted future position $\hat{y_{t}}$ and the closest ground-truth detection $d_{t}^{c}$ at each future timestep, summing over the $T_{pred}$ timesteps as follows:
$d_{t}^{c}=\underset{m\in M}{\operatorname{argmin}}\|\hat{y}_{t}-d_{t}^{m}\|_{2},$ (3)
@@
$\mathcal{L}_{ft}=\mathcal{L}_{W}+\lambda\mathcal{L}_{Reg},$ (6)
@@
Refer to caption Refer to caption Refer to caption Refer to caption